Fowler is expected to pay a dividend of $1.85 one year from today and $2.00 two years from today. The company has a dividend payout ratio of 25 percent and the PE ratio is 19.15 times. If the required return on the company's stock is 12.1 percent, what is the current stock price?
Answer #1

Price at end of 2 years =PE ratio *Dividend year 2/Dividend Payout =19.15*2/(25%) =153.20
Current Stock Price =1.85/(1+12.1%)+2/(1+12.1%)^2+153.20/(1+12.1%)^2=125.15